\begin{thebibliography}{9}                                                                                                %


\bibitem {B1}Bouldin, R., The essential minimum modulus, Indiana Univ. Math.
J. 30 (1981), no. 4, 513--517.

\bibitem {B2}Bouldin R., Distance to invertible linear operators without
separability, Proc. Amer. Math. Soc. 116 (1992), no. 2, 489--497.

\bibitem {BDF}Brown L., Douglas R., and Fillmore P., Unitary equivalence
modulo the compact operators and extensions of C* -algebras, Proc. Conf.
Operator Theory (Dalhousie Univ., Halifax, N. S. 1973), Lecture Notes in Math.
345, Springer, 1973, 58--128.

\bibitem {H}Hadwin, D., Nonseparable approximate equivalence, Trans. Amer.
Math. Soc. 266 (1981), no. 1, 203--231.

\bibitem {KR}Kadison, R. V. and Ringrose, J. R., Fundamentals of the theory of
operator algebras, Vol. II, Advanced theory. Corrected reprint of the 1986
original. Graduate Studies in Mathematics, 16. Amer. Math. Soc., Providence,
RI, 1997.

\bibitem {KS}Kachkovskiy, I. and Safarov, Y., On the distance to normal
elements in C*-algebras of real rank zero, arXiv:1403.2021v1 [math.OA] 9 Mar 2014.

\bibitem {O}Olsen, C. L., Unitary approximation. J. Funct. Anal. 85 (1989),
no. 2, 392--419.
\end{thebibliography}

\documentclass{amsproc}%
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{graphicx}%
\providecommand{\U}[1]{\protect \rule{.1in}{.1in}}
\theoremstyle{plain}

\newtheorem{corollary}{Corollary}

\newtheorem{lemma}{Lemma}

\newtheorem{remark}{Remark}

\newtheorem{theorem}{Theorem}
\numberwithin{equation}{section}
\begin{document}
\title[Lifting Normals]{Lifting Normal Elements in Nonseparable Calkin Algebras}
\author{Ye Zhang}
\address{University of New Hampshire}
\email{yjg2@unh.edu}
\author{Don Hadwin}
\address{University of New Hampshire}
\email{don@unh.edu}
\urladdr{http://euclid.unh.edu/\symbol{126}don}
\author{Yanni Chen}
\address{University of New Hampshire}
\email{yet2@unh.edu}
\thanks{}
\thanks{Supported in part by a grant from the Simons Foundation}
\subjclass[2010]{ Primary 46L05, 46L10 Secondary 47C15, 47L20}
\keywords{Calkin algebra, nonseparable Hilbert space, distance to the normal operators,
countably cofinal cardinal.}

\begin{abstract}
We use the remarkable distance estimate of Ilya Kachkovskiy and Yuri Safarov,
\cite{KS} to show that if $H$ is a nonseparable Hilbert space and
$\mathcal{K}$ is any closed ideal in $B\left(  H\right)  $ that is not the
ideal of compact operators, then any normal element of $B\left(  H\right)
/\mathcal{K}$ can be lifted to a normal element of $B\left(  H\right)  $.

\end{abstract}
\maketitle

\bigskip

Suppose $H$ is a Hilbert space with $\dim H=d\geq \aleph_{0}$. We let $B\left(
H\right)  $ denote the set of all (bounded linear) operators on $H$. For each
cardinal $m$ with $\aleph_{0}\leq m\leq d,$ we let
\[
\mathcal{F}_{m}\left(  H\right)  =\left \{  T\in B\left(  H\right)
:rankT=\dim \overline{T\left(  H\right)  }<m\right \}  ,
\]
and let $\mathcal{K}_{m}\left(  H\right)  $ be the norm closure of
$\mathcal{F}_{m}\left(  H\right)  $. The set $\left \{  \mathcal{K}_{m}\left(
H\right)  :\aleph_{0}\leq m\leq d\right \}  $ is the collection of all nonzero
proper norm-closed ideals of $B\left(  H\right)  $. The quotient C*-algebra
$\mathcal{C}_{m}\left(  H\right)  =B\left(  H\right)  /\mathcal{K}_{m}\left(
H\right)  $ is called the $m$\emph{-Calkin algebra}, and the quotient map
$\pi_{m}:B\left(  H\right)  \rightarrow \mathcal{C}_{m}\left(  H\right)  $ is
called the $m$\emph{-Calkin map.}

The properties of these Calkin algebras depends significantly on the
properties of the cardinal $m$. We say that $m$ is \emph{countably cofinal} if
$m$ is the supremum of a countable collection of smaller cardinals. It was
shown in \cite[Thm. 4.11]{H} that if $m$ is \emph{not} countably cofinal
(which holds exactly when $\mathcal{K}_{m}\left(  H\right)  =\mathcal{F}%
_{m}\left(  H\right)  $), then, for any separable unital C*-subalgebra
$\mathcal{A}$ of $\mathcal{C}_{m}\left(  H\right)  $, there is a unital $\ast
$-homomorphism $\rho:\mathcal{A}\rightarrow B\left(  H\right)  $ such that
$\pi_{m}\circ \rho$ is the identity on $\mathcal{A}$. Hence, in this case, for
any element $t$ of $\mathcal{C}_{m}\left(  H\right)  $ there is a $T\in
B\left(  H\right)  $ so that $\pi_{m}:C^{\ast}\left(  T\right)  \rightarrow
C^{\ast}\left(  t\right)  $ is a $\ast$-isomorphism sending $T$ to $t$. Hence
$t$ lifts to an element $T$ in $B\left(  H\right)  $ that shares all the
properties preserved under $\ast$-isomorphisms, i.e., being normal, subnormal,
hyponormal, isometric, or unitary.

When the cardinal $m$ is countably cofinal, the situation is not so clear cut.
When $m=\aleph_{0}$, the classical result of L. G. Brown, R. G. Douglas, and
P. A. Fillmore \cite{BDF} shows that the Fredholm index is the only
obstruction to lifting normal elements. When $m>\aleph_{0}$, it is still
possible to lift invertibles and unitaries, but the case of normality has
remained open since 1981. In this paper we prove that when $m>\aleph_{0}$ is
countably cofinal, then, for any normal element $t\in \mathcal{C}_{m}\left(
H\right)  $ there is a normal operator $T\in B\left(  H\right)  $ such that
$\pi_{m}\left(  T\right)  =t$.

A key ingredient in our work is the wonderful new theorem (the
\emph{Kachkovskiy-Safarov inequality}) of \cite{KS}, which estimates the
distance $d\left(  a,\mathcal{N}_{f}\left(  \mathcal{A}\right)  \right)  $ of
an element of $a$ in a C*-algebra $\mathcal{A}$ with real rank zero to the set
$\mathcal{N}_{f}\left(  \mathcal{A}\right)  $ of normal operators in
$\mathcal{A}$ with finite spectrum:%
\begin{equation}
d\left(  a,\mathcal{N}_{f}\left(  \mathcal{A}\right)  \right)  \leq C\left(
\left \Vert a^{\ast}a-aa^{\ast}\right \Vert ^{1/2}+d_{1}\left(  a\right)
\right)  , \tag{\#}%
\end{equation}
for some universal constant $C,$ and where
\[
d_{1}\left(  a\right)  =\sup_{\lambda \in \mathbb{C}}dist\left(  a-\lambda
,GL_{0}\left(  \mathcal{A}\right)  \right)  ,
\]
where $GL_{0}\left(  \mathcal{A}\right)  $ is the connected component of $1$
in the group $GL\left(  \mathcal{A}\right)  $ of invertible elements of
$\mathcal{A}$.

When $\mathcal{A}$ is a von Neumann algebra, $GL_{0}\left(  \mathcal{A}%
\right)  =GL\left(  \mathcal{A}\right)  $, so
\[
d_{1}\left(  a\right)  =\sup_{\lambda \in \mathbb{C}}dist\left(  a-\lambda
,GL\left(  \mathcal{A}\right)  \right)
\]
A nice formula for the $\sup dist\left(  b,GL\left(  \mathcal{A}\right)
\right)  $ is given by C. L. Olsen in \cite{O} and more general results appear
in the works of R. Bouldin \cite{B1},\cite{B2}. We will not need these
characterizations here. If $H$ is an infinite-dimensional Hilbert space, then
$H$ is isomorphic to $H\oplus H\oplus \cdots$ , so if $T\in B\left(  H\right)
$, we can identify $T^{\left(  \infty \right)  }=T\oplus T\oplus \cdots$ with an
operator in $B\left(  H\right)  $. Similarly, if $\mathcal{A}$ is an infinite
von Neumann algebra, then there is an orthogonal sequence $\left \{
P_{n}\right \}  $ of projections summing to $1$ so that each $P_{n}$ is
Murray-von Neumann equivalent to $1,$ so if $T\in \mathcal{A}$, we can still
view $T^{\left(  \infty \right)  }$ as an element of $\mathcal{A}$.

\begin{lemma}
\label{T infinity}Suppose $\mathcal{A}$ is an infinite von Neumann algebra
acting on a separable Hilbert space and $T\in \mathcal{A}$. Then

\begin{enumerate}
\item If $\mathcal{A}$ is of type III, then
\[
d_{1}\left(  T\right)  \leq \left \Vert T^{\ast}T-TT^{\ast}\right \Vert ^{1/2},
\]
and%
\[
d\left(  T,\mathcal{N}_{f}\left(  \mathcal{A}\right)  \right)  \leq
2C\left \Vert T^{\ast}T-TT^{\ast}\right \Vert ^{1/2}.
\]

\item If $\mathcal{A}$ is of type I$_{\infty}$ or type II$_{\infty}$, then
\[
d_{1}\left(  T^{\left(  \infty \right)  }\right)  \leq \left \Vert T^{\ast
}T-TT^{\ast}\right \Vert ^{1/2},
\]
and%
\[
d\left(  T^{\left(  \infty \right)  },\mathcal{N}_{f}\right)  \leq2C\left \Vert
T^{\ast}T-TT^{\ast}\right \Vert ^{1/2}.
\]

\end{enumerate}
\end{lemma}

\begin{proof}
Note that $\left \Vert T^{\ast}T-TT^{\ast}\right \Vert ^{1/2}$ is unchanged if
$T$ is replaced with $T-\lambda$, $T^{\left(  \infty \right)  }$ or with
$T^{\ast}$. Hence it will suffice to show $dist\left(  T,GL\left(
\mathcal{A}\right)  \right)  \leq \left \Vert T^{\ast}T-TT^{\ast}\right \Vert
^{1/2}$ in part $\left(  1\right)  $ and $dist\left(  T^{\left(
\infty \right)  },GL\left(  \mathcal{A}\right)  \right)  \leq \left \Vert
T^{\ast}T-TT^{\ast}\right \Vert ^{1/2}$ in part $\left(  2\right)  .$ The parts
involving the distance to $\mathcal{N}_{f}\left(  \mathcal{A}\right)  $ follow
immediately from the Kachkovskiy-Safarov inequality \cite{KS} (see (\#)
above). If both $T^{\ast}T$ and $TT^{\ast}$ are invertible, then $T$ is
invertible and the desired inequalities are trivially true. Since nothing
changes when $T$ is replaced by $T^{\ast}$, there is no harm in assuming that
$T^{\ast}T$ is not invertible. Then, for every $\varepsilon>0$, if we let
$P_{\varepsilon}$ be the spectral projection for $T^{\ast}T$ corresponding to
the interval $\left[  0,\varepsilon \right]  ,$ we have $P_{\varepsilon}\neq0$,%
\[
\left \Vert TP_{\varepsilon}\right \Vert ^{2}=\left \Vert P_{\varepsilon}T^{\ast
}TP_{\varepsilon}\right \Vert \leq \varepsilon
\]
and%
\[
\left \Vert P_{\varepsilon}T\right \Vert ^{2}=\left \Vert P_{\varepsilon}%
TT^{\ast}P_{\varepsilon}\right \Vert \leq \left \Vert P_{\varepsilon}T^{\ast
}TP_{\varepsilon}\right \Vert +\left \Vert T^{\ast}T-TT^{\ast}\right \Vert .
\]
Hence $\left \Vert T-P_{\varepsilon}^{^{\bot}}TP_{\varepsilon}^{^{\bot}%
}\right \Vert \leq \sqrt{\varepsilon}+\sqrt{\varepsilon+\left \Vert T^{\ast
}T-TT^{\ast}\right \Vert }.$

We first consider the case when $\mathcal{A}$ is an infinite \emph{factor}. If
$\mathcal{A}$ is a type $III$ factor, then the projections onto $\ker \left(
P_{\varepsilon}^{^{\bot}}TP_{\varepsilon}^{^{\bot}}\right)  $ and $\ker \left(
P_{\varepsilon}^{^{\bot}}TP_{\varepsilon}^{^{\bot}}\right)  ^{\ast}$ are
nonzero and equivalent, which implies that $P_{\varepsilon}^{^{\bot}%
}TP_{\varepsilon}^{^{\bot}}$ is a norm limit of invertible elements \cite{B2}.
Thus, for every $\varepsilon>0$%
\[
dist\left(  T,GL\left(  \mathcal{A}\right)  \right)  \leq \sqrt{\varepsilon
}+\sqrt{\varepsilon+\left \Vert T^{\ast}T-TT^{\ast}\right \Vert },
\]
which implies%
\[
dist\left(  T,GL\left(  \mathcal{A}\right)  \right)  \leq \left \Vert T^{\ast
}T-TT^{\ast}\right \Vert ^{1/2}.
\]

If $\mathcal{A}$ is a type I$_{\infty}$ or type II$_{\infty}$ factor, then the
projections onto $\ker \left(  P_{\varepsilon}^{^{\bot}}TP_{\varepsilon
}^{^{\bot}}\right)  ^{\left(  \infty \right)  }$ and $\ker \left(  \left(
P_{\varepsilon}^{^{\bot}}TP_{\varepsilon}^{^{\bot}}\right)  ^{\ast}\right)
^{\left(  \infty \right)  }$ are nonzero and have the form $Q^{\left(
\infty \right)  }$ and are equivalent, and we get%
\[
dist\left(  T^{\left(  \infty \right)  },GL\left(  \mathcal{A}\right)  \right)
\leq \left \Vert T^{\ast}T-TT^{\ast}\right \Vert ^{1/2}.
\]
Hence we have proved statements $\left(  1\right)  $ and $\left(  2\right)  $
when $\mathcal{A}$ is an infinite factor von Neumann algebra on a separable
Hilbert space.

For the general case, using the central decomposition \cite{KR}, there is a
direct integral decomposition $\mathcal{A}=\int_{\Omega}^{\oplus}%
\mathcal{A}_{\omega}d\mu \left(  \omega \right)  $, where each $\mathcal{A}%
_{\omega}$ is a factor von Neumann algebra. If $\mathcal{A}$ has type III,
then each $\mathcal{A}_{\omega}$ is a type III factor. If $T=\int_{\Omega
}^{\oplus}T_{\omega}d\mu \left(  \omega \right)  \in \mathcal{A}$ and
$\varepsilon>0$, then, using standard measurable cross-section arguments, it
is easy to measurably choose, for each $\omega \in \Omega$, an invertible
operator $S_{\omega}\in \mathcal{A}_{\omega}$ such that%
\[
\left \Vert T_{\omega}-S_{\omega}\right \Vert \leq \left \Vert T_{\omega}^{\ast
}T_{\omega}-T_{\omega}T_{\omega}^{\ast}\right \Vert ^{1/2}+\varepsilon.
\]
Note that
\[
\left \Vert S_{\omega}\right \Vert \leq \left \Vert T_{\omega}-S_{\omega
}\right \Vert +\left \Vert T_{\omega}\right \Vert \leq \left \Vert T_{\omega}%
^{\ast}T_{\omega}-T_{\omega}T_{\omega}^{\ast}\right \Vert ^{1/2}+\varepsilon
+\left \Vert T_{\omega}\right \Vert \leq
\]%
\[
\left \Vert T^{\ast}T-TT^{\ast}\right \Vert ^{1/2}+\varepsilon+\left \Vert
T\right \Vert \text{ a.e.}\left(  \mu \right)  ,
\]
so $\int_{\Omega}^{\oplus}S_{\omega}d\mu \left(  \omega \right)  $ is defined.
Although each $S_{\omega}$ is invertible, the operator $S=\int_{\Omega
}^{\oplus}S_{\omega}d\mu \left(  \omega \right)  \in \mathcal{A}$ might not be
invertible. However, each $S_{\omega}$ has a polar decomposition $U_{\omega
}\left \vert S_{\omega}\right \vert $ with $U_{\omega}$ unitary. Thus
$S=U\left \vert S\right \vert $ with $U=\int_{\Omega}^{\oplus}U_{\omega}%
d\mu \left(  \omega \right)  $ unitary. Hence, $S$ is the limit of the sequence
$\left \{  U\left(  \left \vert S\right \vert +1/n\right)  \right \}  $ of
invertible operators. Thus,
\[
dist\left(  T,GL\left(  \mathcal{A}\right)  \right)  \leq \left \Vert
T-S\right \Vert \leq \left \Vert T^{\ast}T-TT^{\ast}\right \Vert ^{1/2}%
+\varepsilon \text{.}%
\]
Since $\varepsilon>0$ was arbitrary, we have the desired result. The case when
the $\mathcal{A}_{\omega}$'s are type I$_{\infty}$ or type II$_{\infty}$
factors is handled similarly.
\end{proof}

\begin{corollary}
If $H$ is a Hilbert space and $T\in B\left(  H\right)  $ and $C^{\ast}\left(
T\right)  \cap \mathcal{K}_{\aleph_{0}}\left(  H\right)  =\left \{  0\right \}
,$ then%
\[
d_{1}\left(  T\right)  \leq \left \Vert T^{\ast}T-TT^{\ast}\right \Vert ^{1/2}.
\]

\end{corollary}

\begin{proof}
For every $0\neq A\in C^{\ast}\left(  T\right)  $ we have $rankA\geq \aleph
_{0},$ so $rankA=rankA^{\left(  \infty \right)  }.$ Thus, by \cite[Thm.
3.14]{H}, There is a sequence $\left \{  U_{n}\right \}  $ of unitary operators
in $B\left(  H\right)  $ such that $\left \Vert U_{n}^{\ast}T^{\left(
\infty \right)  }U_{n}-T\right \Vert \rightarrow0$, which means $d_{1}\left(
T\right)  =d_{1}\left(  T^{\left(  \infty \right)  }\right)  $.\bigskip
\end{proof}

We are now ready to prove our main theorem.

\begin{theorem}
Suppose $H$ is an infinite-dimensional Hilbert space and $\aleph_{0}<m\leq \dim
H$ is countably cofinal. If $t\in \mathcal{C}_{m}\left(  H\right)  $ is normal,
then there is a normal operator $T\in B\left(  H\right)  $ such that $\pi
_{m}\left(  T\right)  =t$.
\end{theorem}

\begin{proof}
We first choose an $S\in B\left(  H\right)  $ such that $\pi_{m}\left(
S\right)  =t$. We can write $H$ as a direct sum $H=\sum_{j\in J}^{\oplus}%
H_{j}$ with each dim$H_{j}=\aleph_{0}$ and with $H_{j}$ a reducing subspace
for $S$. Hence, we can write $S=\sum_{j\in J}^{\oplus}S_{j}$ with respect to
this decomposition. Since $S^{\ast}S-SS^{\ast}=\sum_{j\in J}^{\oplus}\left(
S_{j}^{\ast}S_{j}-S_{j}S_{j}^{\ast}\right)  \in \mathcal{K}_{m}\left(
H\right)  $, we see that $E=\left \{  j\in J:S_{j}^{\ast}S_{j}-S_{j}S_{j}%
^{\ast}\neq0\right \}  $ has cardinality at most $m$ and $S_{j}$ is normal when
$j\notin E$. If $CardE<m$, then $T=\sum_{j\in J}^{\oplus}S_{j}\chi_{E}\left(
j\right)  $ is normal and $\pi_{m}\left(  T\right)  =\pi_{m}\left(  S\right)
=t.$ Hence we can assume $J=E,$ so $\dim H=CardE=m$.

It follows from \cite[Cor. 3.11, Thm. 4.6]{H} that there is a unitary operator
$U\in B\left(  H\right)  $ and irreducible operators $A_{1},A_{2},\ldots$ and
cardinals $k_{1},k_{2},\ldots$ such that
\[
U^{\ast}SU-\sum \nolimits_{1\leq n<\infty}^{\oplus}A_{n}^{\left(  k_{n}\right)
}\in \mathcal{K}_{m}\left(  H\right)  .
\]
Hence we can assume that $S=\sum \nolimits_{1\leq n<\infty}^{\oplus}%
A_{n}^{\left(  k_{n}\right)  }$. Since each $A_{n}$ is irreducible, it must
act on a separable Hilbert space. If $k_{n}=m,$ then $A_{n}$ must be normal.
Hence we can write $S=N\oplus \sum \nolimits_{n\in F}^{\oplus}A_{n}^{\left(
k_{n}\right)  }$ , where $k_{n}<m$ whenever $n\in F$. If $F$ is finite,
$\pi_{m}\left(  N\oplus0\right)  =\pi_{m}\left(  S\right)  =t.$ Hence, we can
assume $F=\left \{  n_{1},n_{2},\ldots \right \}  $ with $\aleph_{0}\leq
k_{n_{1}}\leq k_{n_{2}}\leq \cdots<m$ and $m=\sup \left \{  k_{n_{j}}%
:j\in \mathbb{N}\right \}  $. It follows now that
\[
\lim_{j\rightarrow \infty}\left \Vert A_{n_{j}}^{\ast}A_{n_{j}}-A_{n_{j}%
}A_{n_{j}}^{\ast}\right \Vert =0.
\]
However, $A_{j}^{\left(  k_{n_{j}}\right)  }$ is unitarily equivalent to
$\left(  A_{j}^{\left(  \infty \right)  }\right)  ^{\left(  k_{n_{j}}\right)
}$ for each $j,$ which implies by Lemma \ref{T infinity},%
\[
\lim_{j\rightarrow \infty}dist\left(  A_{j}^{\left(  k_{n_{j}}\right)
},\mathcal{N}\right)  =0.
\]
Hence there is a sequence $\left \{  B_{j}\right \}  $ of normal operators such
that
\[
\lim_{j\rightarrow \infty}\left \Vert A_{j}^{\left(  k_{n_{j}}\right)  }%
-B_{j}\right \Vert =0.
\]
Hence $T=N\oplus \sum_{1\leq j<\infty}^{\oplus}B_{j}$ is normal and $\pi
_{m}\left(  T\right)  =\pi_{m}\left(  S\right)  =t$.
\end{proof}

\begin{remark}
If we suppose $m$ is not countably cofinal in the preceding proof, we see that
$CardE$ must be less than $m$ and the proof is complete.
\end{remark}

\end{document}